# Probing Visual Language Priors in VLMs

### v1

@@ 3.2 Question-Image-Answer Generation
As outlined in Criteria Three, we aim for the test answer, $a_{\text{test}}$, to appear highly improbable when considered solely in the context of the question $Q$, yet become a clear and confident choice when paired with the image $I$. This objective presents a significant challenge in finding images $I$, as such images may not follow common knowledge and not naturally exist in the real world or the internet. Recent advancements in generative models offer a solution. By utilizing advanced image generation models such as DALL·E-3 (Ramesh et al., 2021) and Flux, which follow text prompts to synthesize images, we can create images that blend unusual elements. This capability allows us to generate images with the necessary visual coherence to satisfy the specific demands of the question and answer, where we create in a way to disobey language priors. To further ensure the generation of question-image-answers (QIAs) that align with the principles outlined in Section 3.1, our benchmark dataset design incorporates substantial human effort, assisted by state-of-the-art language models such as OpenAI-o1 and Claude-3.5-Sonnet, with additional details provided in Appendix D.1.
Briefly, for each question, we designed three different answers where the first answer $a_{\text{prior}}$ could be inferred solely from the question text. The other two answers were specifically designed against language priors, challenging the model to interpret visual cues despite strong textual priors. During evaluation, we focus on the model’s performance on these second and third answers, i.e., $a_{\text{test}}$. After crafting the initial questions and answers, we leverage GPT-4 to generate text prompts that guide the creation of diverse images at scale. This facilitates the production of varied visual data for the next stage, where high-quality QIA triplets are selected through human review and filtering. Both the text and images are refined as necessary when misalignment or ambiguity are identified. We ensure that the final QIA triplets are clear and easily interpretable by humans, as evidenced by the results of our human evaluation results in Table 2. This iterative process involves several cycles of adjustments to achieve the desired quality. Throughout the process, we encountered two main challenges: (1) generating diverse, out-of-distribution, while reasonable question-answer pairs are difficult, as models like OpenAI-o1 and Claude-3.5-Sonnet tend to converge on a limited set of responses, requiring significant human input to introduce variety; and (2) generating images that against with specific priors using state-of-the-art image generation models is also challenging. In some cases, it was necessary to generate hundreds of images to find just one that accurately matched the question and answer, while avoiding ambiguity.
Table 1: Categorical information.
@@
As a result, we created a total of 300 questions, each paired with three distinct image-answer combinations, resulting in 900 IQA sets. These questions span a broad range of topics, from low-level visual recognition tasks such as texture and shape, to higher-level visual reasoning involving conceptual combinations, hallucinated components, and proverbs. To reinforce the text priors, we also provide a distractor fact before the question, offering additional context for the prior answer $a_{\text{prior}}$ (the one inferred without the image). Some categorical information is shown on the right Table 1. Most of our data are not confined to a single category; instead, they frequently combine multiple categories to challenge even the most advanced VLMs. For instance, in the moon example (6th-row, 2nd-col), we explore shapes and conceptual combinations, such as a round moon paired with a hexagon versus a hexagonal moon. Similarly, the umbrella example (7th-row, 1st-col) incorporates both conceptual combinations and hallucinated components, blending elements like a plunger and a blender. On average, our data span 1.6 categories per question.
@@ 4 Image DPO
Inspired by our benchmark results, we propose Image DPO, a self-improvement method for enhancing VLMs’ visual reasoning, featuring a new objective and a data generation pipeline using VLMs themselves and pre-trained image models.
@@ 4.1 Objective
Figure 2: Illustration of Image DPO. We construct chosen and rejected pairs by corrupting the image with a set of perturbations while keeping the text (questions and answers) unchanged. This setup encourages that the model more focus on the image to differentiate between the chosen and rejected pairs, as both pairs contain identical textual information. The perturbations applied to the images include semantic editing, Gaussian blurring, and pixelization. More details in Appendix C.2.
@@
Existing approaches for VLM self-improvement usually follow the Direct Preference Optimization (DPO) (Rafailov et al., 2024) used in Large Language Models (LLMs), where the model is trained to distinguish between good and bad answers for a fixed image and question. However, this straightforward adaptation is not vision-centered, as the model can still distinguish good and bad answers from the text alone without needing to analyze the image. In contrast, we propose Image DPO, a vision-focused objective that creates good and bad question-image-answer pairs by corrupting the image while keeping the question and answer unchanged (Figure 2).
@@
Mathematically, given the image $I_{w}$, question $Q$ and the corresponding answer $A$, we generate a corrupted image $I_{l}$ by various image editing operations including gaussian blur, pixelation or semantic editing. With the corrupted image $I_{l}$, the triplet of $Q$, $I_{l}$, and $A$ should form a bad question-image-answer pair compared to the original triplet of $Q$, $I_{w}$, and $A$. We train our model to distinguish between good and bad question-image-answer triplets using the objective function:
@@
$L(\theta,\theta_{\text{ref}})=-\mathbb{E}_{Q,I_{w},I_{w},A\sim S}\left[\log\sigma(\alpha\frac{\pi_{\theta}(A\mid Q,I_{w})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{w})}-\alpha\frac{\pi_{\theta}(A\mid Q,I_{l})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{l})})\right]$ (4)
Here, $\theta$ represents the VLM model being trained, $\theta_{\text{ref}}$ is the reference VLM model (typically the previous version of $\theta$), $S$ is the training data consisting of good and bad question-image-answer triplets, $\sigma$ is the sigmoid function, and $\alpha$ is a positive constant.
@@
In Appendix B, we demonstrate that this objective optimizes an upper bound of the RLHF objective (Ouyang et al., 2022), sets it apart from the original DPO objective. Intuitively, since the textual inputs and outputs are identical in both good and bad cases, the gradients of this objective push the model to rely more on the vision branch, driving a shift in gradient direction when processing high-quality images $I_{w}$ compared to corrupted images $I_{l}$, as illustrated in Appendix Figure 9. This behavior encourages the model to focus more on image inputs rather than relying solely on text-based reasoning, thereby enhancing its performance on visually-related tasks.
@@ 4.2 Data Generation
Training VLMs requires a large amount of question-image-answer (QIA) triplets, which are often scarce and expensive to collect. In response, we present a scalable data generation pipeline that creates new QIAs from existing datasets by leveraging VLMs and pre-trained image generation and editing models, as illustrated in Appendix Figure 10. Our pipeline is akin to the tool usage of VLMs/LLMs as prior works (Gupta & Kembhavi, 2022; Parisi et al., 2022; Yao et al., 2022; Sur’is et al., 2023), where pre-trained image models are invoked as callable functions. Given a seed image, VLMs are tasked with simultaneously selecting appropriate functions (e.g., image generation or editing models) and generating corresponding instructions. These instructions are then used to produce new images, in addition to the seed image, as illustrated in Figure 11. The same VLMs are then employed to generate QA pairs for these newly created images. Following, we apply the mentioned three types of image corruptions to the generated images, constructing good-bad QIA pairs (i.e., $I_{w}$ and $I_{l}$). Specifically, we employ Stable Diffusion XL (Podell et al., 2023; Rombach et al., 2022) for image generation, and use Instruct-Pix2Pix (Brooks et al., 2023a), and Grounded-SAM (Rombach et al., 2022; Ren et al., 2024) for image editing. All of our seed images are sourced from existing datasets, including COCO (Lin et al., 2014), Text2VQA (Singh et al., 2019b), and Visual Genome (Krishna et al., 2017). Comprehensive details are provided in Appendix C.
@@ 5 Experiments
We introduce ViLP, a benchmark designed to evaluate VLMs’ visual reasoning capabilities in the presence of strong visual language priors, consisting of 300 carefully crafted questions. Each question is paired with three distinct images and corresponding answers, forming one QIAprior (QIA: question-image-answer) and two QIAtest, resulting in a total of 900 unique QIAs. The QIAprior aligns with common knowledge and language priors, making it easy to infer even from text alone. The QIAtest, on the other hand, challenges the VLM’s visual reasoning ability against these priors.
ViLP  includes two evaluation settings: ViLPF (with distractor facts provided) and ViLPP (pure questions without distractor facts). For each setting, we report the average accuracy on QIAtest as our focused benchmark Score, and the accuracy on QIAprior as Prior. Results are on the Table 2.
Is the QIA easy for humans? We first evaluate our benchmark using human study. Humans achieved nearly 100% accuracy on QIAprior and over 98% accuracy on QIAtest, demonstrating that our question-image-answer combinations are non-ambiguous for humans to answer. Despite QIAtest being designed as out-of-distribution examples, humans were able to distinguish them.
Humans performed slightly better on ViLPF-QIAprior when distractor facts were provided, as they could easily identify that these facts aligned with the correct answers. Moreover, Score on ViLPF-QIAtest was marginally lower when facts were introduced, as the distractor facts added some noise and caused minor confusion, although the impact of this noise is relatively small. These findings are consistent with the design principles of our benchmark.
@@
Figure 4: Before and after removing distactor facts.
Are our QIAs aligned with the learned priors of VLMs? We evaluated GPT-4o’s performance on our questions without providing any images, i.e., GPT-4o (text only) in Table 2. Additionally, we eliminate image-related phrases from the prompt during this text-only inference, such as removing expressions like “as shown in the image". Remarkably, even without any visual content, the model was able to correctly answer 92.33% of QIAprior when distractor facts were included, demonstrating that our questions are highly suggestive for VLMs, allowing them to answer even without visual input. Once removing the distractor facts, the performance decrease to 71.33%, as these facts implicitly pointed to the correct answers in the fact statements. It is noteworthy that some powerful models like Claude-3-Opus and several open-source VLMs fail to outperform GPT-4o (text-only) performance on QIAprior, even when provided with images. For QIAtest, the success rate of GPT-4o (text-only) dropped dramatically to near zero (0% & 0.17%), indicating that QIAtest are far outside the distribution typically encountered by VLMs or LLMs. These observations further validate the design principles of our benchmark: QIAprior can be easily inferred using text alone, while QIAtest is specifically designed to challenge language-based reasoning, requiring reliance on visual cues to arrive at correct answers.
How do VLMs perform on our benchmark? Although our benchmark questions are distinguishable for humans, they are challenging for VLMs. Even the most advanced VLM models like GPT-4o, have a clear performance gap (more than 25%) compared to humans’ performance on QIAtest (Score in ViLPF), indicating the difficulty of these questions for VLMs. GPT-4o achieves above 91% accuracy on QIAprior (Prior), which is quite close to human performance. While its Score on QIAtest is still low (66.17% versus 98.33%), demonstrating the difficulty of these questions for VLMs. We also provide some qualitative examples in Figure 3, showcasing results from both leading commercial and open-source models, including GPT-4o, Claude-3.5-Sonnet, Gemini-1.5-Pro, and Llama-3.2-Vision-90B. They encounter challenges when tackling these cases of our ViLP. Notably, it is encouraging to see that some open-source models achieved over 60% accuracy on ViLPF-QIAtest, with performance nearing that of their commercial counterparts, including Llama-3.2-Vison and Molmo-72B.
Do distractor facts really distract? In ViLPF setting, we add a distractor fact before the question and explicitly instruct the model to answer with one word. Since these facts implicitly suggest incorrect answers for QIAtest, we expected this change to make the questions more suggestive and lower the ViLPFScore, as the distractors would mislead the VLMs.
But is this true in practice? Interestingly, we find that strong models like GPT-4o are not misled by these distractor facts, despite their performance is still significantly lagged behind human accuracy. Not only do these models avoid being misled, but surprisingly, they also leverage the distractor factsto improve their answers, leading to a substantial and consistent performance boost when the facts are included. We illustrate some examples of this behavior in Figure 4. When the distractor fact is provided, GPT-4o performs better on QIAtest compared to when it is not. This result is counterintuitive, as the model performs better when its input prompt includes misleading information. We hypothesize that the distractor fact may help the model infer the focus of the question, thereby narrowing the search space and ultimately leading to higher accuracy.
For models with reasonable instruction-following abilities but limited visual reasoning, like LLaVA-1.5-13B (Liu et al., 2023b), the distractor facts behave as expected, often leading the model to incorrect conclusions based solely on textual information. This leads to a significant performance drop on ViLPF-Score when distractor facts are present, simultaneously boosting performance on ViLPF-Prior, the language-prior answer. For instance, as shown in Figure 4, with including distractor facts, LLaVA-1.5-13B consistently predicts the distractor fact as the answer. However, once the distractors are removed, it can then make the correct prediction.
In contrast, for models with weaker instruction-following abilities, like Cambrian-8B (Tong et al., 2024), the inclusion of distractor facts significantly increases the difficulty of adhering to explicit instructions, such as providing a single-word answer. Specifically, when facts are provided, Cambrian-8B fails to follow the instruction in 62% of the questions, compared to 30% without facts (nearly a 2x increase). Upon manual review of these failures, we find that approximately 59% of the responses that deviate from the instruction are still contextually correct, leading to an adjusted accuracy of 47.92%. A similar issue is observed with LLaVA-OneVision-72B (Li et al., 2024b), which consistently generates sentences with analysis, when explicitly prompted with “please answer this question with one word". This pattern highlights a concerning trend: focusing on improving performance on well-established benchmarks may come at the cost of basic instruction-following abilities, ultimately limiting the practical utility of these models in real-world applications.

### v3

@@ 3.2 Question-Image-Answer Generation
Following Criterion Three, $a_{\text{test}}$ should be highly improbable based on $Q$ alone yet the correct choice when paired with $I$. Since such images do not exist naturally, we use generative models like DALL·E-3 (Ramesh et al., 2021) and Flux to blend unusual elements that override typical language priors. We incorporate substantial human input and leverage advanced LLMs such as OpenAI-o1 and Claude-3.5-Sonnet to ensure alignment with all the criteria. More details are provided in Appendix A.2.
(a) Image-DPO vs. Text-DPO: In the green dashed box, we illustrate Image-DPO, which uses a single Q&A pair paired with multiple corrupted images. In contrast, the purple dashed box presents Text-DPO, involving a single input image paired with multiple, distinct Q&A pairs.
(b) Illustration of Image DPO. We construct chosen and rejected pairs by corrupting the image with a set of perturbations while keeping the Q&A unchanged. Perturbations include semantic editing, Gaussian blurring, and pixelation. The mathematical formulations and implementation details are provided in Appendix B and Appendix C, respectively.
For each question, we design three answers: one $a_{\text{prior}}$ inferred solely from $Q$, and two $a_{\text{test}}$ that defy language priors, requiring visual cues for correctness. We rely on GPT-4 to generate text prompts, produce large-scale images, and then conduct human filtering and refinement. This process faces two main challenges: (1) producing diverse out-of-distribution QA pairs, and (2) synthesizing images that defy specific priors, sometimes necessitating hundreds of samples to find one that accurately matches $Q$ and $a_{\text{test}}$.
Table 1: Category Data.
@@
Ultimately, we curated 300 questions, each paired with three distinct image-answer sets, totaling 900 QIA triplets. These cover a broad range from low-level recognition (texture, shape) to high-level reasoning (conceptual combinations, hallucinated components, proverbs). Table 1 summarizes their categorical distribution, with each question spanning an average of 1.6 categories. To reinforce text priors, we present a distractor fact before the question. Rigorous human review ensures that all final QIA triplets are clear and interpretable, as reflected by our human evaluation results in Table 2.
@@ 4 Image DPO
Inspired by our benchmark, we propose Image DPO, a self-improvement method for enhancing VLMs’ visual reasoning, featuring a new objective and a data generation pipeline using VLMs themselves and pre-trained image models.
@@ 4.1 Objective
Existing approaches for VLM self-improvement follow the way used in DPO paper (Rafailov et al., 2024), where the model is trained to distinguish between good and bad answers for a fixed image and question (Figure 2(a) right). However, this straightforward adaptation may not the best for vision models, as the model sometimes distinguish good and bad answers from the text alone without needing to analyze the image. In contrast, we propose Image DPO, a vision-focused objective that creates good and bad question-image-answer pairs by corrupting the image while keeping the question and answer unchanged (Figure 2(a) left). An example of our synthetic data is illustrated in Figure 2(b).
@@
Formally, given an image $I_{w}$, a question $Q$, and its corresponding answer $A$, we generate a corrupted image $I_{l}$ via image-editing operations, including Gaussian blur, pixelation, or semantic modifications. The triplet $(Q,I_{l},A)$ forms a degraded question-image-answer pair compared to $(Q,I_{w},A)$. We train the model to distinguish between good and bad triplets using the objective below, where $\theta$ is the target VLM, $\theta_{\text{ref}}$ is the reference VLM (typically an earlier version of $\theta$), $S$ is the dataset of good and bad triplets, $\sigma$ is the sigmoid function, and $\alpha$ is a scaling factor.
$L(\theta,\theta_{\text{ref}})$ $=-\mathbb{E}_{Q,I_{w},I_{l},A\sim S}\Bigl{[}$ (4)
@@
$\quad\log\sigma\Bigl{(}\alpha\frac{\pi_{\theta}(A\mid Q,I_{w})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{w})}\;-\;\alpha\frac{\pi_{\theta}(A\mid Q,I_{l})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{l})}\Bigr{)}\Bigr{]}$
@@
Intuitively, since the textual inputs and outputs are identical in both good and bad cases, the gradients of this objective push the model to rely more on the vision branch, driving a shift in gradient direction when processing normal images $I_{w}$ compared to corrupted images $I_{l}$ (Figure 13). This behavior encourages the model to focus more on image inputs rather than relying solely on text-based reasoning, thereby enhancing its performance on visual-related tasks. Our experiments demonstrate Image-DPO objective (Eq. 4) outperforms varies self-improve VLM baselines on ViLP.
Proposition 1.
Let $\mathcal{L}_{\mathrm{RLHF}}(\pi_{\theta},\pi_{\mathrm{ref}};\mathcal{S})$ be the KL-constrained reward maximization objective in Appendix Eq. 7 (Rafailov et al., 2024), where the dataset $\mathcal{S}=\{(Q,A,I_{w},I_{l})\}$ contains good images $I_{w}$ and corrupted images $I_{l}$. Let $\mathcal{L}_{\mathrm{ImageDPO}}(\pi_{\theta},\pi_{\mathrm{ref}};\mathcal{S})$ be the objective from Eq. 4, which compares $\bigl{(}Q,I_{w},A\bigr{)}$ against $\bigl{(}Q,I_{l},A\bigr{)}$. Then for any policy $\pi_{\theta}$ and reference model $\pi_{\mathrm{ref}}$, we have
$\mathcal{L}_{\mathrm{RLHF}}(\pi_{\theta},\pi_{\mathrm{ref}};\mathcal{S})\;\;\leq\;\;\mathcal{L}_{\mathrm{ImageDPO}}(\pi_{\theta},\pi_{\mathrm{ref}};\mathcal{S}).$
Proof Sketch. Following (Rafailov et al., 2024), we express the optimal KL-constrained policy in terms of a latent reward function. Applying a Bradley–Terry preference model to question-image-answer triplets $\bigl{(}Q,I_{w}/I_{l},A\bigr{)}$ and using Jensen’s inequality yields an upper bound whose minimization is equivalent to $\mathcal{L}_{\mathrm{ImageDPO}}(\pi_{\theta},\pi_{\mathrm{ref}};\mathcal{S})$. A full derivation appears in Appendix B.
@@ 4.2 Data Generation
Training VLMs demands large-scale question-image-answer (QIA) triplets, which are often scarce. To address this, we introduce a scalable data generation pipeline (Appendix Figure 14) that repurposes existing image datasets via VLM themselves and image generative models. Given a seed image from COCO (Lin et al., 2014), Text2VQA (Singh et al., 2019b), or Visual Genome (Krishna et al., 2017), VLMs are tasked with simultaneously selecting appropriate functions (e.g., image generation or editing models) and generating corresponding instructions. These instructions are then used to produce new images, in addition to the seed image, as illustrated in Figure 15. The same VLMs are then employed to generate QA pairs for these newly created images. Following, we apply the mentioned three types of image corruptions to the generated images, constructing good bad pars $(Q,I_{w},A)$ and $(Q,I_{l},A)$. Specifically, we employ Stable Diffusion XL (Podell et al., 2023; Rombach et al., 2022) for image generation, and use Instruct-Pix2Pix (Brooks et al., 2023a), and Grounded-SAM (Rombach et al., 2022; Ren et al., 2024) for image editing. Example generated data, prompts, and more details are included in Appendix C.
@@ 5 Experiments
We introduce ViLP,  a benchmark of 300 questions, each with three unique images and corresponding answers (QIAprior and two QIAtest), yielding 900 QIAs. QIAprior aligns with common language priors (easily inferred from text alone), while QIAtest challenges these priors, requiring visual reasoning. ViLP features two evaluation settings: ViLPF (with distractor facts) and ViLPP (pure questions). We report average accuracy on QIAtest (Score) and on QIAprior (Prior) in Table 2.
@@
Is the QIA easy for humans? We begin by evaluating our benchmark through a human study. Participants achieved nearly perfect accuracy on ViLPF-Prior and over 98% on ViLPF-Score and ViLPP-Score, confirming that our question-image-answer combinations are unambiguous for human interpretation. Notably, despite QIAtest being designed as out-of-distribution examples, humans were still able to correctly distinguish them
Humans performed slightly better on ViLPF-Prior when distractor facts were provided, as they could easily identify that these facts aligned with the correct answers. Moreover, ViLPF-Score was marginally lower when facts were introduced, as the distractor facts added some noise and caused minor confusion, although the impact of this noise is relatively small. These findings are consistent with the design principles of our benchmark.
Figure 5: Qualitative results before and after removing distactor facts. GPT-4o and LLaVA-1.5-13B models yield completely opposite behaviors.
Are our QIAs aligned with the learned priors of VLMs? We tested GPT-4o (text only) on our questions (removing all image references). Despite no visual content, it correctly answered 92.33% on ViLPF-Prior. The accuracy drops to 71.33% once distractor facts are removed, showing that these facts significantly guide the answer. For QIAtest, GPT-4o (text only) accuracy nearly falls to 0% (ViLPF-Score & ViLPP-Score), indicating the QIAtest cannot be answered using text alone.
How do VLMs perform on our benchmark? Although our benchmark questions are distinguishable for humans, they are challenging for VLMs. Even the advanced VLM models like GPT-4o, have a clear performance gap (66.17% v.s. 98.33%) compared to humans’ performance on ViLPF-Score, indicating the difficulty of these questions for VLMs. Claude-3.5-Sonnet achieved the best score 70%, while most of the commercial VLMs are below 60%. Figure 3 highlights sample outputs from top commercial and open-source models, including GPT-4o, Claude-3.5-Sonnet, Gemini-1.5-Pro, and Llama-3.2-Vision-90B. They face significant challenges when addressing these cases in our ViLP, whereas humans can arrive at correct answers after consideration. Notably, it is encouraging to see that some open-source models achieved over 60% accuracy on ViLPF-Score, with performance nearing that of their commercial counterparts, including Llama-3.2-Vison and Molmo-72B.
Table 3: Effectiveness of Image-DPO on General VQA benchmarks.
[…]
LLaVA-1.5-7B 29.67 37.67 37.7 43.8 12.7 67.3 31.1 49.1 14.8
LLaVA-1.5-7B + Image-DPO 34.17$\uparrow$ 4.5 39.33$\uparrow$ 1.66 39.79$\uparrow$ 2.09 45.47$\uparrow$ 1.67 14.16$\uparrow$ 1.46 68.45 $\uparrow$ 1.15 32.3$\uparrow$ 1.2 45$\uparrow$ -4.1 12.3$\uparrow$ -2.5
LLaVA-1.5-13B 35.33 41.5 39.6 44.6 14.8 68.9 36.1 48.3 14.1
LLaVA-1.5-13B + Image-DPO 38.17$\uparrow$ 2.84 42.5$\uparrow$ 1 42.68$\uparrow$ 3.08 47.37$\uparrow$ 2.77 17.16$\uparrow$ 2.36 70.36$\uparrow$ 1.46 37.5$\uparrow$ 1.4 42.6$\uparrow$ -5.7 11.6$\uparrow$ -2.5
Cambrian-8B 8.67 32.5 44.6 47.9 19.4 71.5 51.4 14.5 4.7
Cambrian-8B + Image-DPO 20.83 $\uparrow$ 12.16 39.3 $\uparrow$ 6.83 46.5 $\uparrow$ 1.9 50.2 $\uparrow$ 2.3 20 $\uparrow$ 0.6 72$\uparrow$ 0.5 51.7$\uparrow$ 0.3 11.4 $\uparrow$ -3.1 4.4 $\uparrow$ -0.3
Do distractor facts really distract? In ViLPF setting, we add a distractor fact before the question. Since these facts implicitly suggest incorrect answers for QIAtest, we expected this change to make the questions more suggestive and lower the ViLPF-Score, as the distractors would mislead the VLMs. Surprisingly, GPT-4o benefits from distractor facts, improving accuracy on QIAtest. We hypothesize these facts highlight question focus, narrowing the search space. However, weaker models like LLaVA-1.5-13B (Liu et al., 2023b) often get misled by the distractors, hurting their Score but boosting Prior. For instance, as shown in Figure 5, with including distractor facts, LLaVA-1.5-13B consistently predicts the distractor fact as the answer. However, once the distractors are removed, it can then predict correctly.
For bad instruction-following models like Cambrian-8B (Tong et al., 2024), distractor facts significantly hinder adherence to explicit instructions, such as providing single-word answers. With facts, Cambrian-8B fails to follow instructions in 62% of cases, compared to 30% without (a nearly 2x increase). Manual review shows 59% of these failures are contextually correct, yielding an adjusted accuracy of 47.92%. Similarly, LLaVA-OneVision-72B (Li et al., 2024b) often generates detailed analyses despite explicit single-word prompts. This trend highlights a concerning trend: focusing on improving performance on well-established benchmarks may come at the cost of basic instruction-following abilities, ultimately limiting the practical utility of these models in real-world applications.
@@
Comparison of ViLP with other VQA datasets: To highlight the distinctions between ViLP and existing benchmarks, including Winoground (Thrush et al., 2022), Whoops! (Bitton-Guetta et al., 2023a), and HallusionBench (Guan et al., 2023), we conduct a comparative analysis of both their high-level design principles and low-level data formats. This comparison incorporates qualitative and quantitative insights, offering a comprehensive understanding of their differences. Due to space constraints, detailed analyses are provided in Appendix A.3.